OpenStudy (anonymous):

How do you simplify 16 and five fourths?

OpenStudy (tkhunny):

\(16\dfrac{5}4{}\)? This is equivalent to \(16 + \dfrac{5}{4} = \dfrac{64}{4} + \dfrac{5}{4}\) Are we getting anywhere?

OpenStudy (anonymous):

I guess I should be more specific. It's along the lines of simplifying radical expressions to rational.

OpenStudy (phi):

you mean simplify \[ 16^\frac{5}{4} \]? write it as \[16^{\frac{1}{4}\cdot 5}\\ \left( 16^\frac{1}{4}\right)^5 \] 16 has an integer 4th root

OpenStudy (tkhunny):

Yes, you should give the correct problem and you should ALWAYS show your work. So far, you have shown none.